What is the present value of $ 10 comma 000 received a. 12 years from today when the interest rate is 4 % per​ year? b. 20 years from today when the interest rate is 8 % per​ year? c. 6 years from today when the interest rate is 2 % per​ year?
Answer #1

a. Present value = $10,000 / (1.04)12 = $6,245.97

b. Present value = $10,000 / (1.08)20 = $2,145.48

c. Present value = $10,000 / (1.02)6 = $8,879.71